Market 3: Qs = 15 + 7P3 Qd = 25 -2P1 + 3P2 - 4P3
Answer #1

Please give ratings it will be appreciable, for any query please comment, Thank you

Solution,

Market Clearing Condition

Qd = Qs

Market 1

Qs = -5+2P_1\,\,\,\, Qd=15-2P_1-3P_3\\ Qs=Qd\\ -5+2P_1=15-2P_1-3P_3\\ 2P_1+2P_1=15 + 5-3P_3\\ 4P_1=20 - 3P_3 \\P_1=\frac{20 - 3P_3}{4}

Market 2

Qs = 6+2P_2\,\,\,\, Qd=30-3P_2-2P_3\\ Qs=Qd \\ 6+2P_2=30-3P_2-2P_3 \\ 2P_2+3P_2=30-6-3P_3 \\ 5P_2=24 - 2P_3 \\P_2=\frac{24 - 4P_3}{5}

Market 3

Qs = 15+7P_3\,\,\,\, Qd=25-2P_1+3P_2-4P_3 \\ Qs=Qd \\ 15+7P_3=25-2P_1+3P_2-4P_3 \\ 7P_3+4P_3 =25-15 - 2P_1+3P_2 \\ 11P_3=10 - 2P_1 + 3P_2,

\\Since,\,P_1=\frac{20 - 3P_3}{4},\,P_2=\frac{24 - 4P_3}{5} \\11P_3 = 10 - 2\left(\frac{20 - 3P_3}{4}\right) +3\left(\frac{24 - 4P_3}{5}\right) \\11P_3 = 10 - (\frac{20 - 3P_3}{2}) +(\frac{72 - 12P_3}{5}) \\11P_3 = \frac{100 - (100 - 15P_3) +(144 - 24P_3)}{10} \\11P_3 = \frac{100 - 100 + 15P_3 +144 - 24P_3}{10} \\110P_3 =144 - 9P_3

110P3 + 9P3 = 144

P3 = 144/119 = 1.21

\\Since,\,P_3 = 1.21 \\P_1=\frac{20 - 3P_3}{4},\, and \,P_2=\frac{24 - 4P_3}{5} \\P_1 =\frac{20 - 3(1.21)}{4} = \frac{20 - 3.63}{4} = \frac{16.37}{4} = 4.09 \\P_2 = \frac{24 - 4(1.21)}{5} = \frac{24 - 4.84}{5} =\frac{19.16}{5} = 3.83

Price in Market 1 = 4.09

Price in Market 2 = 3.83

Price in Market 3 = 1.21